OpenStudy (anonymous):

find the odds of an event if the probability of the event given to be 2/3

OpenStudy (kropot72):

Odds are the ratio of the probability of an event occurring to that of its not occurring. The probability of the event occurring is 2/3, therefore the probability of the event not occurring is 1 - 2/3 = 1/3. The odds of the event occurring to its not occurring are: \[\frac{2}{3}:\frac{1}{3}\ or\ 2:1\]